# Affective Multimodal Agents with Proactive Knowledge Grounding for Emotionally Aligned Marketing Dialogue

### v1

@@ I Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionised conversational AI, enabling dialogue systems to achieve an unprecedented level of linguistic fluency and contextual understanding [1, 2, 44]. However, despite these remarkable capabilities, most current dialogue agents remain fundamentally reactive, responding to user inputs without proactively guiding conversations toward specific goals [3]. This limitation becomes particularly pronounced in goal-oriented domains such as marketing, where successful interactions require not only understanding user intent but also recognizing emotional states, adapting persuasion strategies, and maintaining long-term engagement [4].
Traditional marketing dialogue systems face several critical challenges that limit their effectiveness in real-world applications. First, they often operate in a purely textual modality, missing crucial emotional cues from voice intonation, facial expressions, and body language that human sales representatives naturally leverage [5]. Second, these systems typically rely on static knowledge bases that cannot adapt to evolving product information, user preferences, or market trends in real-time [6]. Third, existing approaches lack sophisticated emotional intelligence, failing to recognize, understand, and appropriately respond to the complex emotional dynamics that characterize successful sales interactions [7].
The importance of emotional alignment in marketing conversations cannot be overstated. Research in consumer psychology has consistently shown that purchasing decisions are heavily influenced by emotional factors, with emotional connection to a brand or product often outweighing purely rational considerations [8, 9]. Moreover, customers expect increasingly personalized and empathetic interactions from AI systems, particularly in high-involvement purchase scenarios [10]. Current dialogue systems, however, struggle to maintain emotional coherence across extended conversations and often produce responses that feel mechanical or emotionally misaligned with user needs.
@@ II-A Multimodal Dialogue Systems
Recent advances in multimodal dialogue systems have focused on integrating visual and textual information to create more contextually aware conversational agents [11, 12]. Early multimodal dialogue work primarily combined pre-trained vision and language models through simple concatenation or attention mechanisms [13, 14]. More sophisticated approaches have emerged with the development of transformer-based architectures that can jointly process multiple modalities [15, 16].
Recent work has explored various fusion strategies for multimodal dialogue [17]. Early fusion approaches process raw multimodal inputs simultaneously [18], while late fusion methods combine modality-specific representations at higher levels [19]. Intermediate fusion techniques, such as cross-modal attention mechanisms [20], have shown promise in balancing computational efficiency with representation quality.
@@ II-B Affective Computing in Dialogue
Affective computing has emerged as a critical component of human-computer interaction, with particular relevance to dialogue systems [7, 21]. Early emotion recognition approaches focused on single modalities, such as text-based sentiment analysis [22] or speech emotion recognition [23]. More recent work has explored multimodal emotion recognition that combines facial expressions, voice characteristics, and linguistic content [24, 25].
In dialogue contexts, emotion recognition serves multiple purposes: understanding user emotional states, generating emotionally appropriate responses, and maintaining emotional coherence across conversations [26]. Several recent works have attempted to incorporate emotion awareness into dialogue generation. For example, MIME [27] uses emotion recognition to condition response generation, while EmpDialogue [28] focuses on empathetic response generation.
However, the majority of extant affective dialogue systems treat emotion as a secondary consideration rather than a primary driving force for conversation strategy. Prior studies also show that although empathetic or emotion-aware responses can improve user experience, they rarely integrate emotional factors into persuasion-oriented dialogue planning [46, 47]. Existing systems tend to prioritise emotional appropriateness over emotional persuasiveness, despite evidence that emotional expression and affective framing significantly influence persuasive effectiveness [48, 49, 50]. This oversight neglects the pivotal link between affective states and persuasive impact that typifies effective marketing and decision-making interactions.
@@ II-C Knowledge Grounding and Proactive Dialogue
Knowledge grounding in dialogue systems involves integrating external knowledge sources to provide informative and factually accurate responses [6, 29]. Traditional approaches rely on static knowledge bases or retrieval systems that provide relevant facts based on conversation context [30, 31].
Recent work has explored dynamic knowledge integration approaches that can update knowledge representations during conversations [32, 33]. These systems typically use memory networks or attention mechanisms to selectively incorporate relevant knowledge while maintaining conversation coherence.
Proactive dialogue represents a significant departure from traditional reactive systems. Instead of merely responding to user inputs, proactive agents actively guide conversations toward specific objectives [34, 35]. This capability is particularly important in task-oriented domains such as customer service, sales, and education, where the agent must balance user needs with organizational goals.
@@ II-D Marketing Dialogue and Persuasive AI
The intersection of AI and marketing has generated significant research interest, particularly in the development of conversational marketing systems [10, 36]. Early marketing chatbots focused primarily on information provision and basic customer service functions [37]. More sophisticated systems have begun incorporating persuasion techniques derived from social psychology and behavioral economics [38, 39].
Persuasive AI systems face unique challenges related to user trust, transparency, and ethical considerations [40]. Recent work has explored various persuasion strategies in AI systems, including social proof, authority, scarcity, and reciprocity [41, 42]. However, most existing persuasive AI systems operate through simple rule-based approaches rather than sophisticated emotional intelligence.
The ethical implications of persuasive AI have received increasing attention from both researchers and policymakers [43, 45]. Key concerns include user manipulation, privacy protection, and the potential for AI systems to exploit vulnerable populations. Our work addresses these concerns through explicit ethical constraints and transparency mechanisms.
@@ IV-A1 Implementation Details
Our implementation uses PyTorch 1.12 with CUDA 11.7 for GPU acceleration. Training was conducted on 4 NVIDIA A100 GPUs with 40GB memory each. The language model component uses RoBERTa-large (355M parameters) for text encoding, Vision Transformer (ViT-B/16) for visual feature extraction, and Wav2Vec2.0-large for audio processing. The total model contains approximately 1.2B parameters across all components.
@@ IV-A2 Datasets
: A specialized dataset focusing on emotional dynamics in promotional conversations, featuring 5,000 annotated sessions with detailed emotion labels (using a 6-class emotion model: joy, sadness, anger, fear, surprise, disgust) and persuasion success indicators. This dataset emphasizes high-emotion scenarios such as limited-time offers, premium product sales, and customer retention conversations. Sessions average 22.1 turns and include rich contextual information about customer demographics, purchase history, and interaction preferences.
@@ IV-A3 Baseline Methods
: A BERT-based dialogue model enhanced with visual and audio feature fusion through concatenation and cross-modal attention mechanisms [16].
@@
: A recent vision-language model adapted for dialogue generation through fine-tuning on conversational data [12].
@@
: An enhanced version of the empathetic dialogue system with additional marketing-specific training and multimodal capabilities [28].
@@
: A specialized persuasive dialogue system implementing rule-based persuasion techniques from social psychology literature [4].
@@ IV-B Main Results
Statistical significance testing using paired t-tests confirms that all improvements are statistically significant $p<0.001$. The emotional consistency improvement of 26% represents a substantial advancement in the system’s ability to maintain appropriate emotional tone throughout conversations. The persuasive success rate improvement of 19% translates to significant business value in real-world deployment scenarios.
@@ IV-D Qualitative Analysis
Table III shows performance decay with length, dominated by memory retention (–23.9%) and engagement (–20.4%). Affect-aware compression that preserves (i) objections, (ii) explicit commitments, and (iii) affect-trend descriptors recovers a large share of the drop by keeping strategy selection grounded in what the user last cared about. Peaks align with objection and price-sensitivity turns; the model pivots from spec sheets to social proof (reviews) and logistics (return/warranty) before escalating to scarcity cues, mirroring the quantitative gains in Tables VI and VII. In a high-end electronics vignette, AffectMind de-escalates frustration by reframing “spec overload” into lifestyle outcomes, then grounds claims with verified review snippets; baselines persist with technical detail and lose the user.
@@ VI Conclusion and Future Work
The newly introduced datasets (MM-ConvMarket and AffectPromo) provide valuable resources for future research in multimodal marketing dialogue. These datasets capture the complexity of real-world commercial interactions and enable standardized evaluation of emotional AI capabilities.

### v2

@@ I Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionised conversational AI, enabling dialogue systems to achieve an unprecedented level of linguistic fluency and contextual understanding [brown2020language, zhang2025ccma, liang2024comprehensive]. However, despite these remarkable capabilities, most current dialogue agents remain fundamentally reactive, responding to user inputs without proactively guiding conversations toward specific goals [zhang2020dialogpt]. This limitation becomes particularly pronounced in goal-oriented domains such as marketing, where successful interactions require not only understanding user intent but also recognizing emotional states, adapting persuasion strategies, and maintaining long-term engagement [wang2022persuasive].
Traditional marketing dialogue systems face several critical challenges that limit their effectiveness in real-world applications. First, they often operate in a purely textual modality, missing crucial emotional cues from voice intonation, facial expressions, and body language that human sales representatives naturally leverage [poria2017review]. Second, these systems typically rely on static knowledge bases that cannot adapt to evolving product information, user preferences, or market trends in real-time [dinan2018wizard]. Third, existing approaches lack sophisticated emotional intelligence, failing to recognize, understand, and appropriately respond to the complex emotional dynamics that characterize successful sales interactions [picard2000affective].
The importance of emotional alignment in marketing conversations cannot be overstated. Research in consumer psychology has consistently shown that purchasing decisions are heavily influenced by emotional factors, with emotional connection to a brand or product often outweighing purely rational considerations [damasio1994descartes, bechara2005role]. Moreover, customers expect increasingly personalized and empathetic interactions from AI systems, particularly in high-involvement purchase scenarios [kumar2019artificial]. Current dialogue systems, however, struggle to maintain emotional coherence across extended conversations and often produce responses that feel mechanical or emotionally misaligned with user needs.
@@ II-A Multimodal Dialogue Systems
Recent advances in multimodal dialogue systems have focused on integrating visual and textual information to create more contextually aware conversational agents [liang2024comprehensive, liang2022foundations, li2023blip]. Early multimodal dialogue work primarily combined pre-trained vision and language models through simple concatenation or attention mechanisms [antol2015vqa, das2017visual]. More sophisticated approaches have emerged with the development of transformer-based architectures that can jointly process multiple modalities [li2019visualbert, lu2019vilbert].
Recent work has explored various fusion strategies for multimodal dialogue [zeng2025tcstnet]. Early fusion approaches process raw multimodal inputs simultaneously [baltruvsaitis2018multimodal], while late fusion methods combine modality-specific representations at higher levels [zadeh2017tensor]. Intermediate fusion techniques, such as cross-modal attention mechanisms, have shown promise in balancing computational efficiency with representation quality.
@@ II-B Affective Computing in Dialogue
Affective computing has emerged as a critical component of human-computer interaction, with particular relevance to dialogue systems [picard2000affective, calvo2010affect]. Early emotion recognition approaches focused on single modalities, such as text-based sentiment analysis [liu2012sentiment] or speech emotion recognition [el2011survey]. More recent work has explored multimodal emotion recognition that combines facial expressions, voice characteristics, and linguistic content [baltruvsaitis2016openface, zadeh2018multimodal].
In dialogue contexts, emotion recognition serves multiple purposes: understanding user emotional states, generating emotionally appropriate responses, and maintaining emotional coherence across conversations [zhou2018emotional]. Several recent works have attempted to incorporate emotion awareness into dialogue generation. For example, MIME [majumder2020mime] uses emotion recognition to condition response generation, while EmpDialogue [rashkin2019towards] focuses on empathetic response generation.
However, the majority of extant affective dialogue systems treat emotion as a secondary consideration rather than a primary driving force for conversation strategy. Prior studies also show that although empathetic or emotion-aware responses can improve user experience, they rarely integrate emotional factors into persuasion-oriented dialogue planning [samad2022empathetic, shi2021refine]. Existing systems tend to prioritise emotional appropriateness over emotional persuasiveness, despite evidence that emotional expression and affective framing significantly influence persuasive effectiveness [yoshino2018dialogue, gencc2025persuasion, carrasco2024large]. This oversight neglects the pivotal link between affective states and persuasive impact that typifies effective marketing and decision-making interactions.
@@ II-C Knowledge Grounding and Proactive Dialogue
Knowledge grounding in dialogue systems involves integrating external knowledge sources to provide informative and factually accurate responses [dinan2018wizard, ghazvininejad2018knowledge]. Traditional approaches rely on static knowledge bases or retrieval systems that provide relevant facts based on conversation context [moon2019opendialkg, zhou2018commonsense].
Recent work has explored dynamic knowledge integration approaches that can update knowledge representations during conversations [madotto2018mem2seq, wu2019proactive]. These systems typically use memory networks or attention mechanisms to selectively incorporate relevant knowledge while maintaining conversation coherence.
Proactive dialogue represents a significant departure from traditional reactive systems. Instead of merely responding to user inputs, proactive agents actively guide conversations toward specific objectives [tang2019target]. This capability is particularly important in task-oriented domains such as customer service, sales, and education, where the agent must balance user needs with organizational goals.
@@ II-D Marketing Dialogue and Persuasive AI
The intersection of AI and marketing has generated significant research interest, particularly in the development of conversational marketing systems [kumar2019artificial, chung2020chatbot]. Early marketing chatbots focused primarily on information provision and basic customer service functions [folstad2017chatbots]. More sophisticated systems have begun incorporating persuasion techniques derived from social psychology and behavioral economics [fogg2002persuasive, torning2009persuasive].
Persuasive AI systems face unique challenges related to user trust, transparency, and ethical considerations [berdichevsky1999toward]. Recent work has explored various persuasion strategies in AI systems, including social proof, authority, scarcity, and reciprocity [ham2015should, stibe2015advancing]. However, most existing persuasive AI systems operate through simple rule-based approaches rather than sophisticated emotional intelligence.
The ethical implications of persuasive AI have received increasing attention from both researchers and policymakers [ryan2020conversational, metz2021conversational]. Key concerns include user manipulation, privacy protection, and the potential for AI systems to exploit vulnerable populations. Our work addresses these concerns through explicit ethical constraints and transparency mechanisms.
@@ IV-A1 Implementation Details
Our implementation uses PyTorch 1.12 with CUDA 11.7 for GPU acceleration. Training was conducted on 8 HGX H100 node GPUs with 80GB memory each. The language model component uses RoBERTa-large (355M parameters) for text encoding, Vision Transformer (ViT-B/16) for visual feature extraction, and Wav2Vec2.0-large for audio processing. The total model contains approximately 1.2B parameters across all components.
@@ IV-A2 Datasets
: A specialized dataset focusing on emotional dynamics in promotional conversations, featuring 5,000 annotated sessions with detailed emotion labels (using a 6-class emotion model: Positive, Neutral, Negative, Angry, Confused, Excited) and persuasion success indicators. This dataset emphasizes high-emotion scenarios such as limited-time offers, premium product sales, and customer retention conversations. Sessions average 22.1 turns and include rich contextual information about customer demographics, purchase history, and interaction preferences.
@@ IV-A3 Baseline Methods
: A BERT-based dialogue model enhanced with visual and audio feature fusion through concatenation and cross-modal attention mechanisms [lu2019vilbert].
@@
: A recent vision-language model adapted for dialogue generation through fine-tuning on conversational data [li2023blip].
@@
: An enhanced version of the empathetic dialogue system with additional marketing-specific training and multimodal capabilities [rashkin2019towards].
@@
: A specialized persuasive dialogue system implementing rule-based persuasion techniques from social psychology literature [wang2022persuasive].
@@ IV-B Main Results
Statistical significance testing using paired t-tests confirms that all improvements are statistically significant. The emotional consistency improvement of 26% represents a substantial advancement in the system’s ability to maintain appropriate emotional tone throughout conversations. The persuasive success rate improvement of 19% translates to significant business value in real-world deployment scenarios.
@@ IV-D Qualitative Analysis
Table III shows performance decay with length, dominated by memory retention (–23.9%) and engagement (–20.4%). Affect-aware compression that preserves (i) objections, (ii) explicit commitments, and (iii) affect-trend descriptors recovers a large share of the drop by keeping strategy selection grounded in what the user last cared about. Peaks align with objection and price-sensitivity turns; the model pivots from spec sheets to social proof (reviews) and logistics (return/warranty) before escalating to scarcity cues, mirroring the quantitative gains in Tables VI and VII. In a high-end electronics vignette, AffectMind de-escalates frustration by reframing ”spec overload” into lifestyle outcomes, then grounds claims with verified review snippets; baselines persist with technical detail and lose the user.
@@ V-C Limitations and Future Work
Dataset Availability
: The MM-ConvMarket and AffectPromo datasets were collected in collaboration with retail industry partners under standard commercial agreements. Due to the inclusion of real customer interactions and business-sensitive conversion data, the raw datasets are subject to privacy and confidentiality requirements that prevent public release. Interested researchers may contact the corresponding author to explore potential collaboration opportunities.
Reproducibility Constraints
[…]
@@ VI Conclusion and Future Work
The proprietary datasets (MM-ConvMarket and AffectPromo) developed through industry partnerships demonstrate the potential of emotion-grounded dialogue systems in commercial contexts.